# Transforming Indoor Localization: Advanced Transformer Architecture for NLOS Dominated Wireless Environments with Distributed Sensors

### v1

@@ VI-B Performance Analysis of the Proposed Tokenization Methods
The FLOPs complexity and model specifications are greatly influenced by the tokenization approach used. Table I details the Transformer model specifications, with three values in each category being: {$N_{\rm L}$, $D_{\text{emb}}$, $h_{\text{dim}}$}. For instance, {5,12,18}, vanilla Transformer architecture consists of 5 encoder blocks with embedding dimension $D_{\text{emb}}=12$, and hidden dimension $h_{\text{dim}}=18$. The number of attention heads is fixed at $N_{\rm h}=6$, and $\sigma$ is defined as the Rectified Linear Unit (ReLU) for all approaches. For the PBT framework, $W_{\rm h}$ and $W_{\rm w}$ are defined as 8 and 3, respectively, resulting in a total of $N_{\rm tk}=96$ tokens. Each token is represented as a vector of length $N_{\rm st}=24$.
@@
The performances of the three tokenization methods are evaluated by analyzing the Cumulative Distribution Function (CDF) of the 2D positioning error for the device, as illustrated in Fig. 7. The figure presents the CDF results for a large Vanilla Transformer model evaluated across small, medium, and large datasets (from left to right) using all three tokenization methods. It is observed that the proposed SST tokenization method consistently outperforms the other two methods, namely PBT and TST, achieving a significant performance improvement in all three cases. In Fig. 7(c) SST method outperforms other tokenization methods, achieving a 90th percentile ($\Delta_{90}$) error of $0.388$ m. In comparison, PBT and TST report $\Delta_{90}$ errors of $0.659$ m and $0.694$ m, respectively. These findings indicate that SST effectively captures environmental dependencies and multivariate correlations, enhancing model performance, while the other methods fuse information from distinct sensors which leads to poor performance.
@@
It can be observed that as the dataset size increases (from Fig. 7(a) to Fig. 7(c)), the CDF curve shifts to the right, indicating a reduction in positioning error. Specifically, for the $\Delta_{90}$ error, the TST tokenization method demonstrates a percentage improvement of 26.34% when the dataset increases from small to medium, and $40.28$% when increasing from small to large. Similarly, for the PBT tokenization method, the percentage improvement is $24.08$% for a small to medium dataset increase, and $37.77$% for a small to large dataset increase. In comparison, the proposed SST tokenization method exhibits a percentage improvement of $26.47$% as the dataset grows from small to medium, and $40.75$% when transitioning from small to large. These results highlight that while all tokenization methods benefit from larger datasets, the proposed SST achieves superior improvements. Furthermore, the comparatively lower percentage improvements of TST and PBT suggest that their suboptimal tokenization strategies limit the model’s ability to effectively leverage larger datasets, whereas SST demonstrates its capability to harness the additional data for enhanced performance.
@@
Fig. 8 illustrates the $\Delta_{90}$ errors for small, medium, and large models across datasets of varying sizes. As shown in Fig. 8(a), the $\Delta_{90}$ error for all model sizes trained on the small dataset indicates that the proposed SST consistently achieves the lowest positioning errors. A similar trend is observed in Fig. 8(b) and 8(c), where the Transformer models are trained on medium and large datasets. Comparing the $\Delta_{90}$ error of the small, medium, and large models trained on the large dataset, the proposed tokenization method achieves percentage improvements of $76.06$%, $58.53$%, and $44.08$%, respectively, over the TST. Despite having the same computational complexity in terms of FLOPs, the performance gap between the models is notable, emphasizing the effectiveness of the proposed tokenization method.
@@
Considering model complexity, the model trained on a small dataset using SST (see Fig. 8(a)) achieves superior positioning accuracy. Notably, the small model trained with the proposed SST tokenization method not only outperforms models of similar size but also surpasses both medium and large models trained with PBT and TST tokenization techniques. Specifically, the small model using SST achieves a $70.92$% improvement over models with comparable computational complexity. Furthermore, it demonstrates a $46.31$% improvement compared to a model $3.66$ times larger trained with PBT and achieves a $22.38$% improvement over a model $14.1$ times larger trained with PBT. These results highlight that efficient tokenization can achieve higher accuracy with lower computational requirements, as also observed in Fig. 8(b) and 8(c).
(a) Small model.
(b) Medium model.
(c) Large model.
@@
Figure 9: Percentage improvement in 2D positioning errors of the proposed L-SwiGLU Transformer over the Vanilla Transformer using proposed SST tokenization for small, medium, and large models across three datasets at the 50th, 67th, 80th, and 90th percentiles.
@@
VI-C Performance Analysis of the Proposed L-SwiGLU ViT
@@
The proposed L-SwiGLU model architecture shares a similar parameter configuration with the Vanilla Transformer for SST (Table. I), differing primarily in the hidden dimension, denoted as $h^{\text{SwiGLU}}_{\text{dim}}$. For the small, medium, and large models, the values of $h^{\text{SwiGLU}}_{\text{dim}}$ are $54$, $94$, and $231$, respectively. These values are chosen such that the model complexity, in terms of FLOPs, remains equivalent to that of the Vanilla Transformer, ensuring a fair comparison. The number of encoder blocks ($N_{\rm L}$), embedding dimension ($D_{\text{emb}}$), and attention heads ($H=6$) remain identical to those of the Vanilla Transformer, as summarized in Table I.
As the proposed SST tokenization method has consistently demonstrated superior performance across all cases, this subsection focuses exclusively on evaluating the proposed L-SwiGLU Transformer using the SST tokenization.
@@
Fig. 10 presents the CDF of 2D positioning errors for the proposed L-SwiGLU Transformer with SST tokenization and compares it against the Vanilla Transformer utilizing all three tokenization methods: PBT, TST, and SST. The results correspond to the scenario where a large model is trained on a large dataset. Despite maintaining similar computational complexity, the proposed L-SwiGLU Transformer achieves notably better performance, outperforming the Vanilla Transformer. Specifically, the $\Delta_{90}$ positioning error ($\Delta_{90}$) for the Vanilla Transformer is $0.388$ m, whereas the proposed L-SwiGLU Transformer achieves a reduced error of $0.355$ m, demonstrating its superior accuracy and efficiency in large-scale scenarios.
@@
The Fig. 9 demonstrates the percentage improvement in 2D positioning errors achieved by the L-SwiGLU Transformer over the Vanilla Transformer with proposed SST method across small, medium, and large models for three dataset sizes (small, medium, and large) at the 50th, 67th, 80th, and 90th percentiles. It can be seen that the proposed L-SwiGLU Transformer model demonstrates a positive percentage improvement, indicating its superior performance over the Vanilla Transformer. Across all model configurations (Fig. 9(a) to Fig. 9(c)), the large dataset consistently delivers the most significant performance gains, as it provides richer and more diverse training samples, enabling the model to better learn complex patterns and generalize effectively.
@@
For the large dataset, the small model demonstrates a percentage improvement ranging from $10.15$% to $11.18$%, while the medium model shows a decrease to $8.91$% to $7.22$%. For the large model, the improvement ranges from $10.08$% to $8.51$%. The higher percentage improvement observed in the small model is attributed to its previously limited ability to fully utilize the large dataset when trained with the Vanilla Transformer. With the proposed L-SwiGLU Transformer model, the small model is now able to effectively leverage the large dataset, resulting in significant performance gains.
@@
For the small and medium models (Fig. 9(a), and Fig. 9(b)), improvements are moderate for the medium dataset and highest for the large dataset, while the small dataset yields limited gains. For the large model, the performance improvement on the small dataset is minimal at the 50th and 67th percentiles. However, at the ($\Delta_{90}$), the proposed L-SwiGLU Transformer demonstrates superior performance. As the model size increases, the percentage improvement becomes more sensitive to dataset size, reinforcing the need for appropriately scaled datasets to realize the full potential of larger models.
@@
Figure 11: Comparison of $\Delta_{90}$ positioning errors for the L-SwiGLU Transformer across varying model sizes on the large dataset, with the Vanilla Transformer using three tokenization methods.
@@
Fig. 11 presents the $\Delta_{90}$ positioning error for the proposed L-SwiGLU Transformer using SST compared to the Vanilla Transformer across different tokenization methods on the large dataset. The L-SwiGLU Transformer with SST consistently achieves superior performance. Despite having similar computational complexity, the percentage improvement of the L-SwiGLU Transformer over the Vanilla Transformer (both using SST) ranges from $7.22$% to $10.15$%. Compared to the Vanilla Transformer using TST, the percentage improvements are $78.48$%, $61.50$%, and $48.85$% for the small, medium, and large datasets, respectively. Similarly, improvements over the Vanilla Transformer with PBT are $74.37$%, $56.95$%, and $46.13$%. The proposed L-SwiGLU Transformer demonstrates superior performance compared to the Vanilla Transformer models with PBT of significantly larger sizes. Specifically, the small L-SwiGLU Transformer outperforms the medium Vanilla Transformer trained with PBT, which is $3.75$ times larger, achieving a $66.99$% improvement. It also surpasses the large Vanilla Transformer with PBT, which is $14.1$ times larger, with a $46.13$% improvement. When compared to the Vanilla Transformer with SST, the improvements over PBT are $46.31\%$ and $22.38\%$, respectively, as detailed in Subsection VI-B. These results highlight the substantial performance improvements achieved by the proposed SwiGLU, with percent gains significantly increasing from $46.31\%$ to $66.99\%$ and from $22.38\%$ to $46.13\%$. These results highlight the effectiveness of the proposed SST tokenization and the significant performance gains achieved by the L-SwiGLU Transformer.

### v2

@@ VI-B Performance Analysis of the Proposed Tokenization Methods
The FLOPs complexity and model specifications are greatly influenced by the tokenization approach used. Table I details the Transformer model specifications, with three values in each category being: $N_{\rm L}$, $D_{\text{emb}}$, $h_{\text{dim}}$}. For instance, {5,12,18}, vanilla-T architecture consists of 5 encoder blocks with embedding dimension $D_{\text{emb}}=12$, and hidden dimension $h_{\text{dim}}=18$. The number of attention heads is fixed at $N_{\rm h}=6$, and $\sigma$ is defined as the Rectified Linear Unit (ReLU) for all approaches. For the PBT framework, $W_{\rm h}$ and $W_{\rm w}$ are defined as 8 and 3, respectively, resulting in a total of $N_{\rm tk}=96$ tokens. Each token is represented as a vector of length $N_{\rm st}=24$.
@@
The performances of the three tokenization methods are evaluated by analyzing the Cumulative Distribution Function (CDF) of the 2D positioning error for the device, as illustrated in Fig. 7. The figure presents the CDF results for a large Vanilla-T model evaluated across small, medium, and large datasets (from left to right) using all three tokenization methods. It is observed that the proposed SST tokenization method consistently outperforms the other two methods, namely PBT and TST, achieving a significant performance improvement in all three cases. In Fig. 7(c) SST method outperforms other tokenization methods, achieving a 90th percentile ($\Delta_{90}$) error of $0.388$ m. In comparison, PBT and TST report $\Delta_{90}$ errors of $0.659$ m and $0.694$ m, respectively. These findings indicate that SST effectively captures environmental dependencies and multivariate correlations, enhancing model performance, while the other methods fuse information from sensors, which leads to poor performance.
@@
It can be observed that as the dataset size increases (from Fig. 7(a) to Fig. 7(c)), the CDF curve shifts to the left, indicating a reduction in positioning error. Specifically, for the $\Delta_{90}$ error, the TST tokenization method demonstrates a percentage improvement of 26.34% when the dataset increases from small to medium, and 40.2840.28% when increasing from small to large. Similarly, for the PBT tokenization method, the percentage improvement is 24.0824.08% for a small to medium dataset increase, and 37.7737.77% for a small to large dataset increase. In comparison, the proposed SST tokenization method exhibits a percentage improvement of 26.4726.47% as the dataset grows from small to medium, and 40.7540.75% when transitioning from small to large. These results highlight that while all tokenization methods benefit from larger datasets, the proposed SST achieves superior improvements. Furthermore, the comparatively lower percentage improvements of TST and PBT suggest that their suboptimal tokenization strategies limit the model’s ability to effectively leverage larger datasets, whereas SST demonstrates its capability to harness the additional data for enhanced performance.
@@
Fig. 8 illustrates the $\Delta_{90}$ errors for small, medium, and large models across datasets of varying sizes. As shown in Fig. 8(a), the $\Delta_{90}$ error for all model sizes trained on the small dataset indicates that the proposed SST consistently achieves the lowest positioning errors. A similar trend is observed in Fig. 8(b) and 8(c), where the Vanilla-T models are trained on medium and large datasets. Comparing the $\Delta_{90}$ error of the small, medium, and large models trained on the large dataset, the proposed tokenization method achieves percentage improvements of 76.0676.06%, 58.5358.53%, and 44.0844.08%, respectively, over the TST. Despite having the same computational complexity in terms of FLOPs, the performance gap between the models is notable, emphasizing the effectiveness of the proposed tokenization method.
@@
Considering model complexity, the model trained on a small dataset using SST (see Fig. 8(a)) achieves superior positioning accuracy. Notably, the small model trained with the proposed SST tokenization method not only outperforms models of similar size but also surpasses both medium and large models trained with PBT and TST tokenization techniques. Specifically, the small model using SST achieves a 70.9270.92% improvement over models with comparable computational complexity. Furthermore, it demonstrates a 46.3146.31% improvement compared to a model $3.66$ times larger trained with PBT and achieves a 22.3822.38% improvement over a model $14.1$ times larger trained with PBT. These results highlight that efficient tokenization can achieve higher accuracy with lower computational requirements, as also observed in Fig. 8(b) and 8(c).
((a))
((b))
((c))
@@
Figure 10: Percentage improvement in 2D positioning errors of the proposed L-SwiGLU-T over the Vanilla-T using proposed SST tokenization for small, medium, and large models across three datasets at the 50th, 67th, 80th, and 90th percentiles.
@@
VI-C Performance Analysis of the Proposed L-SwiGLU-T
@@
The proposed L-SwiGLU-T model architecture shares a similar parameter configuration with the Vanilla-T for SST (Table. I), differing primarily in the hidden dimension, denoted as $h^{\text{SwiGLU}}_{\text{dim}}$. For the small, medium, and large models, the values of $h^{\text{SwiGLU}}_{\text{dim}}$ are $54$, $94$, and $231$, respectively. These values are chosen such that the model complexity, in terms of FLOPs, remains equivalent to that of the Vanilla-T, ensuring a fair comparison. The number of encoder blocks ($N_{\rm L}$), embedding dimension ($D_{\text{emb}}$), and attention heads ($H=6$) remain identical to those of the Vanilla-T, as summarized in Table I. As the proposed SST tokenization method has consistently demonstrated superior performance across all cases, this subsection focuses exclusively on evaluating L-SwiGLU-T using the SST tokenization.
TABLE II: Performance summary: 2D localization error statistics in meters.
Model Mean Std Dev $\Delta_{90}$
Vanilla-T + PBT 0.3672 0.2152 0.6598
Vanilla-T+ TST 0.3953 0.2309 0.6940
Vanilla-T + SST 0.2203 0.1275 0.3880
L-SwiGLU-T + SST 0.1988 0.1181 0.3554
@@
Fig. 9 presents the CDF of 2D positioning errors for the proposed L-SwiGLU-T with SST tokenization and compares it against the Vanilla-T utilizing all three tokenization methods: PBT, TST, and SST. The results correspond to the scenario where a large model is trained on a large dataset. Despite maintaining similar computational complexity, the proposed L-SwiGLU-T achieves notably better performance, outperforming the Vanilla-T. Specifically, the $\Delta_{90}$ positioning error ($\Delta_{90}$) for the Vanilla-T is $0.388$ m, whereas the proposed L-SwiGLU-T achieves a reduced error of $0.355$ m, demonstrating its superior accuracy and efficiency in large-scale scenarios.
To further strengthen the statistical validity of our results, we report the mean, standard deviation, and 90th percentile of the 2D localization error for each method in Table II. As seen, the proposed L-SwiGLU-T with SST achieves the lowest average error of $0.1988$ m with the smallest standard deviation (Std Dev) of 0.1181 m, indicating consistent and accurate localization. This supports the observed CDF trends and addresses concerns about marginal gains being within the noise range.
@@
The Fig. 10 demonstrates the percentage improvement in 2D positioning errors achieved by the L-SwiGLU-T over the Vanilla-T with the proposed SST method across small, medium, and large models for three dataset sizes (small, medium, and large) at the 50th, 67th, 80th, and 90th percentiles. It can be seen that the proposed L-SwiGLU-T model demonstrates a positive percentage improvement, indicating its superior performance over the Vanilla-T. Across all model configurations (Fig. 10(a) to Fig. 10(c)), the large dataset consistently delivers the most significant performance gains, as it provides richer and more diverse training samples, enabling the model to better learn complex patterns and generalize effectively.
@@
For the large dataset, the small model demonstrates a percentage improvement ranging from 10.1510.15% to 11.1811.18%, while the medium model shows a decrease to 8.918.91% to 7.227.22%. For the large model, the improvement ranges from 10.0810.08% to 8.518.51%. The higher percentage improvement observed in the small model is attributed to its previously limited ability to fully utilize the large dataset when trained with the Vanilla-T. With the proposed L-SwiGLU-T model, the small model is now able to effectively leverage the large dataset, resulting in significant performance gains.
To assess the contribution of each architectural component in L-SwiGLU-T, we performed ablation experiments on the small model trained on the large dataset (Fig. 10(a)). Replacing LayerNorm with RMSNorm improved training stability and efficiency. Removing the class token reduced complexity and yielded a 3.4% accuracy gain. Eliminating positional embeddings avoided bias and improved accuracy by 4%, while introducing the SwiGLU feedforward block contributed an additional 2.6% improvement. These results highlight the effectiveness of each modification.
@@
For the small and medium models (Fig. 10(a), and Fig. 10(b)), improvements are moderate for the medium dataset and highest for the large dataset, while the small dataset yields limited gains. For the large model, the performance improvement on the small dataset is minimal at the 50th and 67th percentiles. However, at the ($\Delta_{90}$), the proposed L-SwiGLU-T demonstrates superior performance. As model size increases, improvement becomes more sensitive to dataset size, reinforcing the need for appropriately scaled datasets to realize the full potential of larger models.
@@
Figure 11: Comparison of $\Delta_{90}$ positioning errors for the L-SwiGLU-T across varying model sizes on the large dataset, with the Vanilla-T using three tokenization methods.
@@
Fig. 11 presents the $\Delta_{90}$ positioning error for the proposed L-SwiGLU-T using SST compared to the Vanilla-T across different tokenization methods on the large dataset. The L-SwiGLU-T with SST consistently achieves superior performance. Despite having similar computational complexity, the percentage improvement of the L-SwiGLU-T over the Vanilla-T (both using SST) ranges from 7.227.22% to 10.1510.15%. Compared to the Vanilla-T using TST, the percentage improvements are 78.4878.48%, 61.5061.50%, and 48.8548.85% for the small, medium, and large datasets, respectively. Similarly, improvements over the Vanilla-T with PBT are 74.3774.37%, 56.9556.95%, and 46.1346.13%. The proposed L-SwiGLU-T demonstrates superior performance compared to the Vanilla-T models with PBT of significantly larger sizes. Specifically, the small L-SwiGLU-T outperforms the medium Vanilla-T trained with PBT, which is $3.75$ times larger, achieving a 66.9966.99% improvement. It also surpasses the large Vanilla-T with PBT, which is $14.1$ times larger, with a 46.1346.13% improvement. When compared to the Vanilla-T with SST, the improvements over PBT are $46.31\%$ and $22.38\%$, respectively, as detailed in Subsection VI-B. These results highlight the substantial performance improvements achieved by the proposed SwiGLU, with percent gains significantly increasing from $46.31\%$ to $66.99\%$ and from $22.38\%$ to $46.13\%$. These results highlight the effectiveness of the proposed SST tokenization and the significant performance gains achieved by the L-SwiGLU-T.
VI-D Attention Map Analysis of Proposed L-SwiGLU-T Model
To interpret how the proposed L-SwiGLU-T model utilizes information from distributed sensors under NLOS conditions, we visualize the average attention weights across sensors from selected Transformer layers. Since each layer contains multiple attention heads that capture different representation aspects, we compute the mean attention scores across all six heads to obtain a more interpretable view.
Fig. 12(a) and Fig. 12(b) show the average attention weights from the first and last transformer layers for the small model on the small dataset, respectively, aggregated over six heads. It is evident that both early and deep layers assign significantly higher attention to Sensor 13 and Sensor 14, while suppressing contributions from less informative sensors such as Sensor 1. To further support this behavior, we examine the PDPs of some sensors in Fig. 10. The absence of early, high-power taps and the dispersion of energy across mid-to-late taps across all sensors is a strong indicator of the NLOS nature of the environment. Sensors 13 and 14 exhibit strong peaks at higher taps/delay-bins (e.g., taps 30–45), indicating dominant reflected paths arriving after significant delay. Sensor 1 shows a relatively delayed and diffuse profile, with no dominant early peak, consistent with deep NLOS or heavy obstruction. The attention mechanism clearly learns to assign higher weight to sensors with higher signal quality and lower multipath delay spread, reinforcing the model’s capacity to dynamically prioritize informative spatial observations. This validates the model’s interpretability and effectiveness in leveraging distributed sensor data under multipath-rich indoor environments.
((a))
((b))
((c))
Figure 12: Visualization of attention and signal characteristics under an NLOS scenario. (a)–(b) show average attention weights across sensors from the first and last Transformer layers, respectively, while (c) presents the corresponding PDPs for selected sensors.
VI-E Comparison of L-SwiGLU-T with Lightweight Transformer Baselines
To benchmark efficiency and performance, we compare L-SwiGLU-T with recent state-of-the-art lightweight Transformer architectures. The One Wide Feedforward is All You Need (OWFF) Transformer (OWFF-T) model [1] reduces redundancy by removing separate FFNs per encoder layer and instead employs a single, wide FFN shared across all layers. In contrast, DELIGHT-T (Deep and Light-weight Transformer) [2] introduces a highly compact architecture that employs Group Linear Transformations to capture local representations. To incorporate global context, it applies feature shuffling across groups, analogous to channel shuffling in CNNs.
Fig. 13 plots the CDF of 2D localization errors for the proposed L-SwiGLU-T, OWFF-T, and DELIGHT-T models on the small dataset. All models are constrained to 4.5M FLOPs for fair comparison. The proposed L-SwiGLU-T demonstrates superior performance, achieving a $\Delta_{90}$ error of $0.7974$ m, compared to $0.9316$ m for DELIGHT-T and $0.9562$ m for OWFF-T. This performance gain is attributed to its effective balance between computational efficiency and parameter utilization, leveraging gated activations and structural sparsity to enhance representational capacity without increasing model complexity.
Figure 13: CDF of 2D localization errors for L-SwiGLU-T, OWFF-T, and DELIGHT-T using small model on the small dataset.
VI-F Performance Comparison of L-SwiGLU-T Model Sizes on Real-World Dataset
To demonstrate the practical effectiveness of the proposed L-SwiGLU-T model, we evaluated it on the Ultra-Dense MaMIMO CSI Dataset [nr6k-8r78-21], which provides real-world channel measurements collected using a multi-antenna testbed. The dataset was recorded at a center frequency of 2.61 GHz with a 20 MHz bandwidth and $N=100$ subcarriers. Among the available antenna configurations, we focused on the distributed antenna array (DA) topology, where $N_{S}=8$ sensor nodes are placed at distinct locations surrounding the area of interest. Each sensor node is equipped with $A=8$ receiving antennas, forming a spatially distributed sensing setup. The collected frequency-domain CSI was transformed into TD CIRs using (3) with $N=128$, followed by computation of the PDP via (4). Only the first $N_{\text{ts}}=128$ delay taps were retained for evaluation.
For real-world evaluation using the SST framework, the number of tokens is $N_{\rm tk}=N_{\rm S}=8$, with each token representing one of the eight distributed sensors, and each token having a feature dimension of $N_{\rm st}=N_{\textrm{ts}}=128$. While our simulation studies considered $N_{\rm S}=18$ sensors (see Section IV-C), the proposed L-SwiGLU-T model is designed to support varying sensor counts. Only the number and embedding of input tokens need to be adapted to the deployment; the core model architecture remains unchanged. This demonstrates the flexibility of our SST-based formulation in handling diverse sensor configurations. Furthermore, spatial differences in sensor placement are implicitly learned through the attention mechanism, and the L-SwiGLU-T’s ability to process variable-length sequences enables generalization to dynamic sensor layouts, even without retraining. Another observation is that in this dataset, the number of tokens is reduced to $N_{\rm tk}=N_{\rm S}=8$, compared to 18 in the simulation setup. Since Transformer complexity scales quadratically with the number of tokens, this results in a substantial reduction in computational cost. Specifically, the FLOPs ($\mathcal{B}$) for the small, medium, and large models decrease from 4.5M to 1.8M, 16.5M to 7M, and 63.5M to 27.5M, respectively, highlighting the efficiency of the proposed model in real-world scenarios with fewer sensors.
The real-world dataset contains approximately 250K samples; however, for this study, we utilize a subset of 40K samples, which we refer to as the large dataset in the context of this work. The CDF of the 2D localization error for the small, medium, and large versions of the proposed L-SwiGLU-T is shown in Fig. 14. As model capacity increases, the error consistently decreases, highlighting the model’s ability to capture complex spatial features more effectively. Specifically, the 90th percentile errors are $0.0890$ m for the small model, $0.0714$ m for the medium model, and $0.0483$ m for the large model.
Figure 14: CDF of 2D localization error for small, medium, and large L-SwiGLU-T models on real world dataset.
To benchmark against prior work, we compare our model’s performance with the CNN approach implemented by the authors of the dataset in [de2020csi]. Their CNN model achieved a mean absolute error (MAE) of $0.0823$ m, trained on the entire dataset of 250K samples. In contrast, our proposed L-SwiGLU-T model was trained on a significantly smaller subset of only 40K samples, yet it consistently outperformed the CNN-based based across all model sizes. This highlights the efficiency and generalization capability of the Transformer-based architecture, even with limited data. The results are summarized in Table III.
TABLE III: MAE comparison: Proposed L-SwiGLU-T vs. CNN baseline [de2020csi]
Model MAE (m)
CNN (Baseline, 250K samples) [de2020csi] 0.0823
L-SwiGLU-T Small (40K samples) 0.0482
L-SwiGLU-T Medium (40K samples) 0.0387
L-SwiGLU-T Large (40K samples) 0.0267
These results demonstrate that even under constrained data regimes, the proposed model effectively captures spatiotemporal features with higher precision than convolutional baselines, benefiting from attention mechanisms that model long-range dependencies across distributed sensors.
The proposed SST-based L-SwiGLU-T model is inherently flexible and not restricted to fixed input dimensions such as the number of sensors or PDP length. It readily accommodates different input sizes through token-level adaptation without requiring architectural modifications. As demonstrated in both simulation (with 18 sensors at 3.5 GHz/100 MHz) and real-world evaluation (with 8 sensors at 2.61 GHz/20 MHz), the same model generalizes effectively across diverse sensor configurations and frequency bands, underscoring its robustness for practical deployment. Multi-floor localization and cross-domain adaptation can be effectively achieved using transfer learning, where a pre-trained model can be fine-tuned with minimal data to adapt to new floors or environments.